OpenStudy (anonymous):

I Have THREE question. 1. One day, a person went to horse racing area, Instead of counting the number of human and horses, he instead counted 74 heads and 196 legs. Yet he knew the number of humans and horses there. How did he do it, and how many humans and horses are there? 2. If 1/2x +1/2(1/2x + 1/2(1/2x +1/2(1/2x + ... = y, then x =? 3.*What number shows up most often when you roll 10 dice?

OpenStudy (anonymous):

some random questions here

OpenStudy (anonymous):

Yes I Know, My teacher everyday gives 3 pop quiz questions all different . I got these. They dont make any sense to me help!

OpenStudy (anonymous):

i have no idea what you are studying, or how you are supposed to know this, but the answer to the last question is 35

OpenStudy (anonymous):

thank you. and there POP QUIZ QUESTIONS. I Geuss what we learn in class is what we grow off on this .

OpenStudy (anonymous):

if you have two dice the lowest number you can roll is 2, the highest is 12, half way between them is 7 because \[\frac{2+12}{2}=7\]

OpenStudy (anonymous):

similarly if i you have ten dice the lowest number you can roll is 10, the highest is 60 and half way between them is \[\frac{10+60}{2}=35\] and that will be the most frequent number

OpenStudy (anonymous):

second one is \[y=x\]

OpenStudy (mertsj):

Let h = number of horses Let p = number of people h+p = 74 4h+2p = 196 horses = 24, people = 50

OpenStudy (anonymous):

one way which may or may not make sense is to expand this a as a geometric series and see that you have \[\frac{1}{2}x+\frac{1}{2^2}x+\frac{1}{2^3}x+...=(\frac{1}{2}+\frac{1}{2^2}+\frac{1}{2^3}+...)x=1x\]

OpenStudy (anonymous):

second way is to write \[\frac{x}{2}+\frac{y}{2}=y\] so \[\frac{x}{2}=\frac{y}{2}\] and therefore \[x=y\]

OpenStudy (lilg132):

1) where y is human z is horses total legs = y(2) + z(4) = 196 total heads = y + z = 74 2y = 196 - 4z y = 98 - 2z 98 - 2z + z = 74 98 - 74 = 2z - z 24 = Z y + 24 = 74 y = 74 - 24 y = 50 so there were 50 human heads and 24 horses heads which is the same as 50 humans and 24 horses and with this we also know there were 100 human legs and 96 horses legs dont forget to click good answer :)